OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

@03225186213

OpenStudy (anonymous):

slope of given line is 3 parallel lines have same slope so slope=3 equ y-2.5=3(x-(-10)) y-2.5=3x+30 y=3x+32.5

OpenStudy (anonymous):

thanks :D